# Roadmap for electronic structure, anharmonicity, and electron-phonon calculations in locally disordered inorganic and hybrid halide perovskites

### v1

@@ IV.1 CsPbI3
We first describe our results for cubic CsPbI3 without considering thermal effects. Our DFT-PBEsol calculations for the reference (monomorphous) structure reveal a semimetallic behavior, characterized by a spurious exchange in orbital character between the conduction and valence band extrema [Fig. 12(a)]. We thus assign a negative value to the band gap of $-0.05$ eV, as reported in Table 1. Accounting for positional polymorphism resolves the orbital exchange problem [Fig. 12(b)] and yields a band gap widening of 0.65 eV relative to the monomorphous structure. This band gap increase is the largest among all compounds studied in this work, consistent with the largest energy lowering reported for CsPbI3 in Fig. 2(b). Going beyond DFT and using the PBE0 functional to improve the description of electronic correlations, we obtain a band gap of 1.15 eV for the monomorphous structure. This value is notably lower than the experimental value of 1.78 eV at 650 K Sutton et al. (2018). However, for the locally disordered network, the theoretical value improves significantly, yielding 1.84 eV, showing agreement with the experimental value.
@@ IV.3 CsSnI3
In this section, we discuss our results for cubic CsSnI3. Similar to CsPbI3, our DFT-PBEsol calculations yield a semimetallic behavior for monomorphous CsSnI3, due to the spurious exchange in orbital character between the band extrema. We thus assign a negative value to the band gap of $-0.27$ eV, as reported in Table 1. Taking into account positional polymorphism addresses the band inversion issue and results in a 0.56 eV widening of the band gap compared to the monomorphous structure. By using the PBE0 functional for improved treatment of electronic correlations, we obtain a band gap of 0.67 eV for the monomorphous structure. This value is significantly lower than the experimental value of 1.35 eV at 500 K Kontos et al. (2018). Accounting for local disorder yields a PBE0 value of 1.30 eV, which matches the experimental result.
@@ IV.4 CsPbCl3
[…]
@@ IV.5 MAPbI3
In Fig. 9(m), we report the temperature variation of the hole and electron effective masses in the range 315 – 450 K. Using both ZG reference and ZG polymorphous networks, we obtain a linear variation of the effective masses with temperature. Our analysis for the slopes yields $\lambda^{T}_{{\rm h,ref}}=2.7\times 10^{-4}$ and $\lambda^{T}_{{\rm e,ref}}=1.6\times 10^{-4}$, $\lambda^{T}_{{\rm h,poly}}=4.6\times 10^{-4}$, and $\lambda^{T}_{{\rm e,poly}}=1.0\times 10^{-4}$ $m_{0}/{\rm K}$, showing that the presence of local disorder increase/decrease the slopes $\lambda^{T}_{{\rm h}}$/$\lambda^{T}_{{\rm e}}$ by 70%/38%. As for the band gap renormalization, using a single ZG polymorphous configuration obtained from the MSS structure results in excellent agreement with Boltzmann-weighted averages over 10 polymorphous configurations. We remark that in all cases electron-phonon coupling dominates the thermal induced effective masses enhancements contributing, for example, 97% to $\lambda^{T}_{{\rm h,poly}}$ and 99% to $\lambda^{T}_{{\rm e,poly}}$. The reduced effective mass computed for the ZG polymorphous structure varies linearly from $\mu_{\rm poly}(315~{}{\rm K})=0.096$ to $\mu_{\rm poly}(450~{}{\rm K})=0.110$.
@@ IV.6 MAPbBr3
Figure 9(n) shows temperature-dependent hole and electron effective masses in the range 275 – 450 K calculated using ZG reference (open squares) and ZG polymorphous (red discs) networks. Unlike the hole effective mass of 0.15 $m_{0}$ at 300 K obtained for the reference structures, our calculated value of 0.26 $m_{0}$ for the polymorphous structures at 300 K is in excellent agreement with angle-resolved photoemission measurements at room temperature, which report a value of $0.25\pm 0.05$ $m_{0}$u et al. (2019). Importantly, even when using a single ZG polymorphous configuration obtained from the MSS structure [open circles in Fig. 9(n)], the results compare well with our values obtained from Boltzmann-weighted averages over 10 polymorphous configurations. Linear fits to the data give slopes of $\lambda^{T}_{{\rm h,ref}}=2.5\times 10^{-4}$ and $\lambda^{T}_{{\rm e,ref}}=2.0\times 10^{-4}$, $\lambda^{T}_{{\rm h,poly}}=1.5\times 10^{-4}$, and $\lambda^{T}_{{\rm e,poly}}=0.9\times 10^{-4}$ $m_{0}/{\rm K}$, showing that local disorder decreases the slopes by 40 % and 55 %, close to the values obtained for CsPbBr3. We note that in our calculations, electron-phonon coupling dominates by 85%–92% the thermal-induced effective masses enhancements for both the reference and polymorphous structures. The reduced effective mass obtained for polymorphous cubic MAPbBr3 varies linearly from $\mu_{\rm poly}(275~{}{\rm K})=0.120$ to $\mu_{\rm poly}(450~{}{\rm K})=0.130$.
@@ IV.7 MASnI3
In this section, we discuss our results for cubic MASnI3. Similar to CsPbI3 and CsSnI3, our DFT-PBEsol calculations yield a semimetallic behavior for the reference structures of MASnI3 and a spurious exchange in orbital character between the band extrema. We thus assign a negative value to the band gap of $-0.09$ eV, as reported in Table 1. The band inversion issue is addressed when taking into account positional polymorphism which results in a band gap widening of 0.37 eV. Our PBE0 calculations for the reference structures give a band gap of 0.88 eV, which underestimates the experimental value of 1.31 eV at 300 K Parrott et al. (2016). Employing locally disordered networks results in a band gap opening to 1.30 eV, which is in agreement with experiment and compares well with the $GW$ value of 1.1 eV reported for tetragonal MASnI3 Umari et al. (2014).
@@ IV.8 MAPbCl3
We next investigate the renormalization of the band gap and effective masses due to electron-phonon coupling and thermal expansion. Figure 9(l) shows the band gap of cubic MAPbCl3 in the temperature range 200–400 K obtained using 10 ZG polymorphous (red discs) and 10 ZG reference (open squares) structures for each temperature. We also include temperature-dependent band gaps computed with the ZG polymorphous configuration of the MSS structure (open cirles). Experimental data (gray triangles) are from Ref. [Hsu et al., 2019]. To enable direct comparison with experiment, the DFT data are rigidly shifted by $\Delta_{s}=1.73$ eV. This value matches the PBE0 correction of 1.64 eV obtained for the polymorphous structures (see Table 1). Our data for the ZG reference structures show a temperature-induced band gap widening, with a temperature coefficient of $dE_{\rm g}/dT=6.2\times 10^{-4}$ eV/K. This value significantly overestimates the experimentally derived coefficient of $-0.9\times 10^{-4}$ eV/K, which indicates that the band gap remains nearly constant with temperature in the measured range. As illustrated in Figs. 10(a) and (b), both $dE_{\rm g}/dT|_{\rm EPC}$ and $dE_{\rm g}/dT|_{\rm TE}$ computed for the reference structures lead to band gap widening, with electron-phonon coupling accounting for 48%. Notably, this is the only case among Cs-based and MA-based compounds where thermal expansion is the leading contribution to the thermal induced band gap renormalization. When ZG polymorphous networks are employed, electron-phonon coupling results in band gap narrowing, uniquely among all compounds studied, while thermal expansion continues to induce a band gap widening, though its contribution is reduced. The total temperature coefficient of the band gap for the polymorphous structures is $dE_{\rm g}/dT=0.9\times 10^{-4}$ eV/K, which closely matches the experimental value, as shown in Fig. 10(c). Overall, we attribute the nearly constant variation of the band gap with temperature to the opposing effects of electron-phonon coupling and thermal expansion. Unlike for the other MA-based compounds, a single ZG polymorphous configuration derived from the MSS structure does not provide a sufficiently accurate approximation for computing temperature-dependent band gaps, yielding a temperature coefficient of $3.8\times 10^{-4}$ eV/K.
@@ IV.9 FAPbI3
We next present our results for cubic FAPbI3. DFT-PBEsol calculations for the reference structure reveal a semimetallic behavior with a band gap of 0.10 eV. However, unlike CsPbI3 [Figs. 12(a) and (b)], our charge density plots [Figs. 12(f) and (g)] do not display any spurious exchange in orbital character between the conduction and valence band extrema. Accounting for positional polymorphism leads to a relatively small band gap widening of 0.24 eV, consistent with the relatively small degree of polymorphism in FA-based compounds (Fig. 2). Going beyond DFT, PBE0 calculations give a band gap of 1.33 eV for the reference structures. Although this underestimates the experimental value of 1.53 eV at 300 K Chen et al. (2020), it remains in reasonable agreement. Employing polymorphous networks results in a band gap of 1.59 eV, aligning better with experiment and again highlighting the importance of local disorder in accurate band gap predictions. Our result also compares well with $GW$ calculations for tetragonal FAPbI3, which report a band gap of 1.71 eV Muhammad et al. (2022). Previous $GW$ calculations, based on idealized ferroelectric monomorphous structures that are not representative of the actual physical system yielding an artificial Rashba splitting, report a band gap of 1.47 eV and 1.48 eV Bokdam et al. (2016); Muhammad et al. (2020). We emphasize that band structures of polymorphous structures do not exhibit a Rashba band splitting of the doubly degenerate band extrema Zacharias et al. (2025, 2023a) as they reflect, on average, the system’s centrosymmetricity.
@@ IV.10 FAPbBr3
Our calculated DFT-PBEsol effective masses for the reference structure are $m^{*}_{\rm h,ref}=0.10$ and $m^{*}_{\rm e,ref}=0.10$. Local disorder enhances the effective masses, resulting to $m^{*}_{\rm h,poly}=0.16$ and $m^{*}_{\rm e,poly}=0.15$, and thus to $\lambda_{\rm h}=0.6$ and $\lambda_{\rm e}=0.5$. These enhancements increase the reduced effective mass of cubic FAPbBr3 from $\mu_{\rm ref}=0.050$ to $\mu_{\rm poly}=0.077$ improving the comparison with the experimental values for orthorhombic and tetragonal phases of 0.117 and 0.13 Galkowski et al. (2016). To further improve consistency with experimental values, we apply $GW$ enhancements Wang et al. (2020b) of 72% and 93% to the DFT hole and electron effective masses, respectively, following the approach used for FAPbI3333. This results in a reduced effective mass of $\mu_{{\rm poly}+GW}=0.141$.
@@ V.1 Electronic structure
Improved treatment of electronic correlations using PBE0 calculations combined with polymorphous structures yields overall good agreement with experimental band gap values, as shown in Table 1. The calculated band gaps exhibit an average absolute deviation of 0.11 eV from experimental values. Including thermal induced corrections [$\Delta E_{\rm g,poly}(T)$] further improves agreement in some cases, but increases the discrepancy in others, resulting in a slightly higher average absolute deviation of 0.12 eV. Moreover, our PBE0 values compare well with $GW$ calculations for the tetragonal and orthorhombic phases, which incorporate the effect of octahedral tilting. In future work, it would be interesting to explore how the combination of state-of-the-art $GW$ calculations using cubic polymorphous structures and thermal induced corrections compares with experiments. We emphasize that polymorphism is also essential for improving agreement with experimental values of the reduced effective masses, as shown in Table 1. In this case, $GW$ calculations are also expected to play an important role in achieving better quantitative accuracy. We observe that for the reduced effective masses of Pb-based compounds, the agreement with experiment decreases from Cs-based to MA-based, and further to FA-based compounds. For example, our values underestimate experimental data for CsPbI3, MAPbI3, and FAPbI3 by 18%, 31%, and 48%, respectively. This trend aligns with the fact that compounds with larger cations exhibit smaller degree of polymorphism, and thus stronger electronic correlations.
V.2 LO-TO splitting

### v2

@@ IV.1 CsPbI3
We first describe our results for cubic CsPbI3 without considering thermal effects. Our DFT-PBEsol calculations for the reference (monomorphous) structure reveal a semimetallic behavior, characterized by a spurious exchange in orbital character between the conduction and valence band extrema [Fig. 12(a)]. We thus assign a negative value to the band gap of 0.05-0.05 eV, as reported in Table 1. Accounting for positional polymorphism resolves the orbital exchange problem [Fig. 12(b)] and yields a band gap widening of 0.65 eV relative to the monomorphous structure. This band gap increase is the largest among all compounds studied in this work, consistent with the largest energy lowering reported for CsPbI3 in Fig. 2(b). Going beyond DFT and using the PBE0 functional to improve the description of electronic correlations, we obtain a band gap of 1.15 eV for the monomorphous structure. This value is notably lower than the experimental value of 1.78 eV at 650 K Sutton et al. (2018). However, for the locally disordered network, the theoretical value improves significantly, yielding 1.84 eV, showing agreement with the experimental value.
@@ IV.3 CsSnI3
In this section, we discuss our results for cubic CsSnI3. Similar to CsPbI3, our DFT-PBEsol calculations yield a semimetallic behavior for monomorphous CsSnI3, due to the spurious exchange in orbital character between the band extrema. We thus assign a negative value to the band gap of 0.27-0.27 eV, as reported in Table 1. Taking into account positional polymorphism addresses the band inversion issue and results in a 0.56 eV widening of the band gap compared to the monomorphous structure. By using the PBE0 functional for improved treatment of electronic correlations, we obtain a band gap of 0.67 eV for the monomorphous structure. This value is significantly lower than the experimental value of 1.35 eV at 500 K Kontos et al. (2018). Accounting for local disorder yields a PBE0 value of 1.30 eV, which matches the experimental result.
@@ IV.4 CsPbCl3
The effective masses calculated for the monomorphous structure are $m^{*}_{\rm h,ref}=0.12$ and $m^{*}_{\rm e,ref}=0.13$. Using locally disordered structures, the effective masses increase to $m^{*}_{\rm h,poly}=0.29$ and $m^{*}_{\rm e,poly}=0.31$ which correspond to effective mass enhancement factors of $\lambda_{\rm h}=1.42$ and $\lambda_{\rm e}=1.38$. The corrections arizing from polymorphism improve considerably the agreement with the experimental reduced effective mass of 0.202± $02\,\pm\,0.01$ at $T=2$ K Baranowski et al. (2020), increasing the calculated value for CsPbCl3 from $\mu_{\rm ref}=0.062$ to $\mu_{\rm poly}=0.150$. Despite this important improvement, our computed $\mu_{\rm poly}$ still underestimates experiment by 25%, which we attribute to neglecting quasiparticle and excitonic effects Filip et al. (2021); Lang et al. (2015) from our calculations.
@@ IV.5 MAPbI3
In Fig. 9(m), we report the temperature variation of the hole and electron effective masses in the range 315 – 450 K. Using both ZG reference and ZG polymorphous networks, we obtain a linear variation of the effective masses with temperature. Our analysis for the slopes yields $\lambda^{T}_{{\rm h,ref}}=2.7\times 10^{-4}$ and $\lambda^{T}_{{\rm e,ref}}=1.6\times 10^{-4}$, $\lambda^{T}_{{\rm h,poly}}=4.6\times 10^{-4}$, and $\lambda^{T}_{{\rm e,poly}}=1.0\times 10^{-4}$ $m_{0}/{\rm K}$, showing that the presence of local disorder increase/decrease the slopes $\lambda^{T}_{{\rm h}}$$\lambda^{T}_{{\rm e}}$ by 70%/38%. As for the band gap renormalization, using a single ZG polymorphous configuration obtained from the MSS structure results in excellent agreement with Boltzmann-weighted averages over 10 polymorphous configurations. We remark that in all cases electron-phonon coupling dominates the thermal induced effective masses enhancements contributing, for example, 97% to $\lambda^{T}_{{\rm h,poly}}$ and 99% to $\lambda^{T}_{{\rm e,poly}}$. The reduced effective mass computed for the ZG polymorphous structure varies linearly from $\mu_{\rm poly}(315~{\rm K})=0.096$ to $\mu_{\rm poly}(450~{\rm K})=0.110$.
@@ IV.6 MAPbBr3
Figure 9(n) shows temperature-dependent hole and electron effective masses in the range 275 – 450 K calculated using ZG reference (open squares) and ZG polymorphous (red discs) networks. Unlike the hole effective mass of 0.15 $m_{0}$ at 300 K obtained for the reference structures, our calculated value of 0.26 $m_{0}$ for the polymorphous structures at 300 K is in excellent agreement with angle-resolved photoemission measurements at room temperature, which report a value of $0.25\pm 0.05$ $m_{0}$Zu et al. (2019). Importantly, even when using a single ZG polymorphous configuration obtained from the MSS structure [open circles in Fig. 9(n)], the results compare well with our values obtained from Boltzmann-weighted averages over 10 polymorphous configurations, i.e. when including configurational entropy effects. Linear fits to the data give slopes of $\lambda^{T}_{{\rm h,ref}}=2.5\times 10^{-4}$ and $\lambda^{T}_{{\rm e,ref}}=2.0\times 10^{-4}$, $\lambda^{T}_{{\rm h,poly}}=1.5\times 10^{-4}$, and $\lambda^{T}_{{\rm e,poly}}=0.9\times 10^{-4}$ $m_{0}/{\rm K}$, showing that local disorder decreases the slopes by 40 % and 55 %, close to the values obtained for CsPbBr3. We note that in our calculations, electron-phonon coupling dominates by 85%–92% the thermal-induced effective masses enhancements for both the reference and polymorphous structures. The reduced effective mass obtained for polymorphous cubic MAPbBr3 varies linearly from $\mu_{\rm poly}(275~{\rm K})=0.120$ to $\mu_{\rm poly}(450~{\rm K})=0.130$.
@@ IV.7 MASnI3
In this section, we discuss our results for cubic MASnI3. Similar to CsPbI3 and CsSnI3, our DFT-PBEsol calculations yield a semimetallic behavior for the reference structures of MASnI3 and a spurious exchange in orbital character between the band extrema. We thus assign a negative value to the band gap of 0.09-0.09 eV, as reported in Table 1. The band inversion issue is addressed when taking into account positional polymorphism which results in a band gap widening of 0.37 eV. Our PBE0 calculations for the reference structures give a band gap of 0.88 eV, which underestimates the experimental value of 1.31 eV at 300 K Parrott et al. (2016). Employing locally disordered networks results in a band gap opening to 1.30 eV, which is in agreement with experiment and compares well with the $GW$ value of 1.1 eV reported for tetragonal MASnI3 Umari et al. (2014).
@@ IV.8 MAPbCl3
We next investigate the renormalization of the band gap and effective masses due to electron-phonon coupling and thermal expansion. Figure 9(l) shows the band gap of cubic MAPbCl3 in the temperature range 200–400 K obtained using 10 ZG polymorphous (red discs) and 10 ZG reference (open squares) structures for each temperature. We also include temperature-dependent band gaps computed with the ZG polymorphous configuration of the MSS structure (open cirles). Experimental data (gray triangles) are from Ref. [Hsu et al., 2019]. To enable direct comparison with experiment, the DFT data are rigidly shifted by $\Delta_{s}=1.73$ eV. This value matches the PBE0 correction of 1.64 eV obtained for the polymorphous structures (see Table 1). Our data for the ZG reference structures show a temperature-induced band gap widening, with a temperature coefficient of $dE_{\rm g}/dT=6.2\times 10^{-4}$ eV/K. This value significantly overestimates the experimentally derived coefficient of $0.9\times 10^{-4}$ eV/K, which indicates that the band gap remains nearly constant with temperature in the measured range. As illustrated in Figs. 10(a) and (b), both $dE_{\rm g}/dT|_{\rm EPC}$ and $dE_{\rm g}/dT|_{\rm TE}$ computed for the reference structures lead to band gap widening, with electron-phonon coupling accounting for 48%. Notably, this is the only case among Cs-based and MA-based compounds where thermal expansion is the leading contribution to the thermal induced band gap renormalization. When ZG polymorphous networks are employed, electron-phonon coupling results in band gap narrowing, uniquely among all compounds studied, while thermal expansion continues to induce a band gap widening, though its contribution is reduced. The total temperature coefficient of the band gap for the polymorphous structures is $dE_{\rm g}/dT=0.9\times 10^{-4}$ eV/K, which closely matches the experimental value, as shown in Fig. 10(c). Overall, we attribute the nearly constant variation of the band gap with temperature to the opposing effects of electron-phonon coupling and thermal expansion. Unlike for the other MA-based compounds, a single ZG polymorphous configuration derived from the MSS structure does not provide a sufficiently accurate approximation for computing temperature-dependent band gaps, yielding a temperature coefficient of $3.8\times 10^{-4}$ eV/K.
@@ IV.9 FAPbI3
We next present our results for cubic FAPbI3. DFT-PBEsol calculations for the reference structure reveal a semimetallic behavior with a band gap of 0.10 eV. However, unlike CsPbI3 [Figs. 12(a) and (b)], our charge density plots [Figs. 12(f) and (g)] do not display any spurious exchange in orbital character between the conduction and valence band extrema. Accounting for positional polymorphism leads to a relatively small band gap widening of 0.24 eV, consistent with the relatively small degree of polymorphism in FA-based compounds (Fig. 2). Going beyond DFT, PBE0 calculations give a band gap of 1.33 eV for the reference structures. Although this underestimates the experimental value of 1.53 eV at 300 K Chen et al. (2020), it remains in reasonable agreement. Employing polymorphous networks results in a band gap of 1.59 eV, aligning better with experiment and again highlighting the importance of local disorder in accurate band gap predictions. Our result also compares well with $GW$ calculations for tetragonal FAPbI3, which report a band gap of 1.71 eV Muhammad et al. (2022). Previous $GW$ calculations, based on idealized ferroelectric monomorphous structures that are not representative of the actual physical system yielding an artificial Rashba splitting, report a band gap of 1.47 eV and 1.48 eV Bokdam et al. (2016); Muhammad et al. (2020). We emphasize that band structures of polymorphous structures do not exhibit a Rashba band splitting of the doubly degenerate band extrema Zacharias et al. (2026, 2023a) as they reflect, on average, the system’s centrosymmetricity.
@@ IV.10 FAPbBr3
Our calculated DFT-PBEsol effective masses for the reference structure are $m^{*}_{\rm h,ref}=0.10$ and $m^{*}_{\rm e,ref}=0.10$. Local disorder enhances the effective masses, resulting to $m^{*}_{\rm h,poly}=0.16$ and $m^{*}_{\rm e,poly}=0.15$, and thus to $\lambda_{\rm h}=0.6$ and $\lambda_{\rm e}=0.5$. These enhancements increase the reduced effective mass of cubic FAPbBr3 from $\mu_{\rm ref}=0.050$ to $\mu_{\rm poly}=0.077$ improving the comparison with the experimental values for orthorhombic and tetragonal phases of 0.117 and 0.13 Galkowski et al. (2016). To further improve consistency with experimental values, we apply $GW$ enhancements Wang et al. (2020b) of 72% and 93% to the DFT hole and electron effective masses, respectively, following the approach used for FAPbI33. This results in a reduced effective mass of $\mu_{{\rm poly}+GW}=0.141$.
@@ V.1 Electronic structure
Improved treatment of electronic correlations using PBE0 calculations combined with polymorphous structures yields overall good agreement with experimental band gap values, as shown in Table 1. The calculated band gaps exhibit an average absolute deviation of 0.11 eV from experimental values. Including thermal induced corrections $\Delta E_{\rm g,poly}(T)$] further improves agreement in some cases, but increases the discrepancy in others, resulting in a slightly higher average absolute deviation of 0.12 eV. Moreover, our PBE0 values compare well with $GW$ calculations for the tetragonal and orthorhombic phases, which incorporate the effect of octahedral tilting. In future work, it would be interesting to explore how the combination of state-of-the-art $GW$ calculations using cubic polymorphous structures and thermal induced corrections compares with experiments. We emphasize that polymorphism is also essential for improving agreement with experimental values of the reduced effective masses, as shown in Table 1. In this case, $GW$ calculations are also expected to play an important role in achieving better quantitative accuracy. We observe that for the reduced effective masses of Pb-based compounds, the agreement with experiment decreases from Cs-based to MA-based, and further to FA-based compounds. For example, our values underestimate experimental data for CsPbI3, MAPbI3, and FAPbI3 by 18%, 31%, and 48%, respectively. This trend aligns with the fact that compounds with larger cations exhibit smaller degree of polymorphism, and thus stronger electronic correlations.
V.2 Band gap across different polymorphs
As discussed in Secs. IV and V.1, the band gaps obtained for the polymorphous cubic structures agree well with prior calculations on lower-symmetry polymorphs that explicitly incorporate octahedral tilts, i.e. the tetragonal and orthorhombic phases. Considering the monomorphous phases of CsPbBr3, our calculations yield a DFT-PBEsol band gap of $E_{\mathrm{g}}=0.83$, $0.69$, and $0.24$ eV for the orthorhombic (Pnma), tetragonal (P4/mbm), and cubic ($\mathrm{Pm}\bar{3}\mathrm{m}$) phases, respectively. Assuming a monomorphous picture this trend is reasonable: larger octahedral tilts in the orthorhombic phase widen the gap relative to the tetragonal phase (a0a0c±; no apical tilting) and, ultimately, to the untilted cubic phase. However, this ordering contradicts the measured band gap sequence in CsPbBr3 Mannino et al. (2020), where $E_{\mathrm{g}}(\mathrm{Pm}\bar{3}\mathrm{m})>E_{\mathrm{g}}(\mathrm{P4/mbm})>E_{\mathrm{g}}(\mathrm{Pnma})$. The corresponding calculated band gaps using polymorphous structures are: $E_{\mathrm{g}}=0.83$, $0.86$, and $0.85$ eV for the orthorhombic, tetragonal, and cubic phases, respectively. This nearly recovers the expected trend (orthorhombic << tetragonal $\simeq$ cubic), compressing the spread to 0.03 eV. As shown in Ref. [Zacharias et al., 2023a], adopting polymorphous rather than monomorphous structures and including thermal effects is crucial to capture the continuous increase of the band gap with temperature across different phases, avoiding spurious discontinuities at the phase boundaries. This suggests that reproducing the correct band gap ordering requires explicitly accounting for local structural disorder and finite-temperature effects, i.e. thermal lattice expansion and electron-phonon coupling.
A similar trend holds for MAPbBr3 Mannino et al. (2020). Using monomorphous orthorhombic, reference tetragonal, and reference cubic structures of MAPbBr3, our DFT-PBEsol calculations yield $E_{\mathrm{g}}=0.83$, $0.46$, and $0.39$ eV, respectively. In contrast, polymorphous tetragonal and cubic structures that capture local disorder give an almost phase-independent gap of $E_{\mathrm{g}}\simeq 0.83$ eV. Interestingly, representing the finite-temperature band gap with a ZG orthorhombic MAPbBr3 snapshot already provides a better description of the band gap at high temperatures than using ZG reference cubic cells Zacharias et al. (2026).
For MAPbI3, the picture differs slightly. Monomorphous orthorhombic, reference tetragonal, and reference cubic structures give $E_{\mathrm{g}}=0.61$, $0.29$, and $0.03$ eV, respectively, having a spread of 0.58 eV. Employing polymorphous tetragonal and cubic structures the corresponding band gaps are $E_{\mathrm{g}}=0.61$, $0.51$, and $0.52$ eV, reducing the spread to 0.1 eV; this residual offset reflects the gap drop at the orthorhombic to tetragonal transition, not observed in MAPbBr3 Mannino et al. (2020). This result is consistent with experimental observations for MAPbI3 Milot et al. (2015); Dar et al. (2016) which report a band gap drop at the orthorhombic to tetragonal phase transition of about 0.1 eV and a nearly smooth evolution from the tetragonal to cubic phase transition.
V.3 LO-TO splitting